OpenStudy (anonymous):

Connor bought 30 shares of eBay at $20.95. A year later, he bought 20 more shares at $26.75. He later sold all of his shares at the price of $35.25. Connor's broker charges him 3% of each share as a transaction fee. After all of his transactions, how much money did Connor make? A. $511.21 B. $493.37 C. $525.11 D. $515.53

OpenStudy (anonymous):

A. (35.25*50-20.95*30-26.75*20) - ((20.95*30)+(26.75*20*)+(35.25*50))*.03 = 599-87.78 = $511.22